# SIC-Free Rate-Splitting Multiple Access: Constellation-Constrained Optimization and Application to Large-Scale Systems

### v1

@@ V Optimization for Large-Scale Systems
To solve linear precoding design problems for large-scale systems, the inherent challenge is that the grown size of the design variable, i.e. precoding matrix, leads to growing computational complexity. For problems considering finite constellations, the even more fatal issue is the exponential growth in the cardinality of observed constellations at each user. It can be observed from (5) and (6) that the computational complexity of evaluating the CC sum-rate grows exponentially with respect to the number of received signal sources and the cardinality of the constellations adopted by those input signals. An estimate of computational complexity of the sum-rate of a MU-MIMO system with different number of users and constellation cardinality is given in Table II in terms of floating point operations (FLOPs). Practically, evaluating scenarios with more than 12 bits transmitted over all the streams becomes difficult.
@@
2 $1.18\times 10^{8}$ $2.77\times 10^{10}$ $4.05\times 10^{12}$
@@
To enable precoding designs and evaluation for large-scale systems with practical computational burden and acceptable reduction in performance, we introduce the following techniques, namely user grouping and null space projection. The idea is to partition the co-scheduled users into groups, enable one common stream for each group, and constrain the precoders such that inter-group interference is eliminated. Figure 2 illustrates a system in which $K$ users (assuming that $K$ is even) are partitioned into $\frac{K}{2}$ groups, and each group contains two users. The elimination of inter-group interference is illustrated since signals carrying the common and private streams intended for a particular user do not propagate to other groups, although intra-group interference, i.e. interference caused by the private stream intended for a co-group user, potentially exists. In contrast to Table II, an estimate of computational complexity with user grouping and interference nulling is given by Table II, where the growth in FLOPs is roughly linear w.r.t. the number of users. To address the growth in the dimension of precoding matrices w.r.t. the number of antennas, we also utilize a technique to reduce the dimension of the design variables to further accelerate optimization algorithms.
@@ V-A User Grouping
Figure 2: Illustration of user grouping.
It is intuitive that, with optimal precoding matrix, a user would suffer less interference from streams intended for co-scheduled users which have dissimilar channels. Hence, we heuristically group users with similar channels to limit the potential performance degradation. For the purpose of limiting the computational burden, the number of users within each group should be limited. We consider a special case where each group contains two users. The proposed user grouping algorithm is described in Alg. 5666We assume $K$ to be even for clarity, while extension to odd numbers is straightforward. For example, one can assume that one group contain only one user, while all other groups contains two users., where $q(\mathbf{h}_{m},\mathbf{h}_{n})$ is a predefined channel similarity metric. We use the angle of channels for our analysis, i.e.
@@
though alternative metrics might provide superior results or serve different purposes.
@@
2:  Compute $\mathbf{Q}\in\mathcal{C}^{K\times K}$, where$[\mathbf{Q}]_{m,n}=q(\mathbf{h}_{m},\mathbf{h}_{n})$, $\forall m<n,m\in\mathcal{K},n\in\mathcal{K}$, and$[\mathbf{Q}]_{m,n}=-1$, $\forall m\geq n,m\in\mathcal{K},n\in\mathcal{K}$.
@@
It is evident that Alg. 5 can be extended to handle cases with more than two users per group.
@@ V-B Null Space Projection and Dimensionality Reduction
$\mathbf{p}_{\mathcal{K}_{i},\text{c}}=\mathbf{F}_{\mathcal{K}_{i}}\mathbf{G}_{\mathcal{K}_{i}}\mathbf{v}_{\mathcal{K}_{i},\text{c}},$ (31)
@@
where $\mathbf{F}_{\mathcal{K}_{i}}$ projects the subsequent vector onto the null space of channels of users outside $\mathcal{K}_{i}$, such that inter-group interference is eliminated. We denote and express the inter-group channels to users in $\mathcal{K}_{i}$ as
@@
where the columns of $\mathbf{H}_{\mathcal{K}_{j}}$ are the channel vectors of users in group $\mathcal{K}_{j}$. It is readily seen that inter-group interference nulling can be achieved by having
@@
Apart from interference nulling, an additional benefit of $\mathbf{F}_{\mathcal{K}_{i}}$ is that the dimensionality of the subsequent design variable is reduced from $N_{\text{T}}$ to $N_{\text{T}}-K+2$, i.e. the dimensionality of the null space of inter-group channels. We introduce $\mathbf{G}_{\mathcal{K}_{i}}$ to further reduce the dimensionality of the design variable by extending the low-dimensional subspace property (LSP) introduced by [41] to finite constellation scenarios as follows.
@@
With inter-group interference nulling, the LSP implies that optimal $\mathbf{G}_{\mathcal{K}_{i}}\mathbf{v}_{\mathcal{K}_{i},\text{c}}$ and $\mathbf{G}_{\mathcal{K}_{i}}\mathbf{v}_{k}$ lies in the column space of $\mathbf{F}_{\mathcal{K}_{i}}^{H}\mathbf{H}_{\mathcal{K}_{j}}$. This allows dimensionality reduction in the design variable from $N_{\text{T}}-K+2$ to $2$ while ensures the optimal solution can still be achieved. Let
$\mathbf{F}_{\mathcal{K}_{i}}^{H}\mathbf{H}_{\mathcal{K}_{j}}=\mathbf{U}_{\mathcal{K}_{i}}\mathbf{\Sigma}\mathbf{V}_{\mathcal{K}_{i}}^{H}$ (35)
be compact singular value decomposition of $\mathbf{F}_{\mathcal{K}_{i}}^{H}\mathbf{H}_{\mathcal{K}_{j}}$. We restrict $\mathbf{G}_{\mathcal{K}_{i}}\mathbf{v}_{k}$ to be in the subspace by choosing [^7] To achieve the same purpose, one can also have $\mathbf{G}_{\mathcal{K}_{i}}=\mathbf{F}_{\mathcal{K}_{i}}^{H}\mathbf{H}_{\mathcal{K}_{j}}$. However, it was observed numerically that the proposed $\mathbf{G}_{\mathcal{K}_{i}}$ leads to better convergence when optimizing $\mathbf{v}_{k}$. One potential reason is that the proposed $\mathbf{G}_{\mathcal{K}_{i}}$ guarantees that $\|\mathbf{p}_{k}\|^{2}=\|\mathbf{v}_{k}\|^{2}$ and $\|\mathbf{p}_{\mathcal{K}_{i},\text{c}}\|^{2}=\|\mathbf{v}_{\mathcal{K}_{i},\text{c}}\|^{2}$.
@@
With the first and second stages of precoding in closed form by (34) and (36). The third stage, that is, $\mathbf{v}_{k}$ and $\mathbf{v}_{\mathcal{K}_{i},\text{c}}$, is designed in a similar manner as in Section IV. Note that, as a result of inter-group interference nullification, inter-group interference does not need to be considered when computing the achievable rates and their gradients for $\mathcal{P}_{1}$ and $\mathcal{P}_{3}$. For the MMF problem, as each group posses a common stream, the design of common rate allocation, i.e. Algorithm 3, is carried out for each group separately.
@@ VI-A Channel Model
where $\mathbf{h}_{k}^{\text{LoS}}(\theta_{k}^{\text{az}},\theta_{k}^{\text{el}})$, $\mathbf{h}_{k}^{\text{NLoS}}$ and $\kappa_{k}$ denote the Line-of-Sight (LoS) channel component, non-LoS channel component and and the Rician factor. $\mathbf{h}_{k}^{\text{LoS}}$ is characterized by the array geometry and the direction of departure in azimuth and elevation, denoted by $\theta_{k}^{\text{az}}$ and $\theta_{k}^{\text{el}}$ respectively. $\mathbf{h}_{k}^{\text{NLoS}}$ is modeled by Rayleigh fading.
In this work, we aim to model channels in practical scenarios where users are in high-density demand areas. For small-scale systems, a uniform linear array is assumed along the $y$ axis with $\theta_{k}^{\text{el}}=0,\forall k\in\mathcal{K}$ since elevation does not affect the channels. For large-scale systems, a uniform rectangular array along the $y$-$z$ plane is assumed. It is assumed that $\kappa_{1}=...=\kappa_{K}=\kappa$.
@@ VI-B Adaptive Modulation
The advantage of RSMA lies in its flexibility in adjusting the power and rates between common and private streams. To fully exploit this advantage under finite constellations, we allow RSMA to dynamically adjust the constellations. In particular, we enable RSMA to maximize the objective function by exhaustively searching for constellations from a pre-defined dictionary of transmission modes. To reduce the searching complexity, private streams are limited to using the same constellations in each transmission mode. To ensure fairness in decoding complexity, different modes within the same dictionary are designed to produce the same maximum throughput, denoted by $R_{\text{max}}$. According to the above design principles, the four transmission mode dictionaries used in the simulations are documented in Table VI - VI in Appendix C. Note that with Mode 1 in all the examples, RSMA is equivalent to SDMA, since the precoding optimization will allocate zero power to the common stream when the common stream constellation is empty. For large-scale systems, we assume that all the groups use the same transmission mode, which is searched from one of the pre-defined dictionaries. In the numerical results, where necessary, $K$ and $R_{\text{max}}$ are provided in the legends or captions to identify the transmission mode dictionary in use.

### v2

@@ V Optimization for Large-Scale Systems
To solve linear precoding design problems for large-scale systems, the inherent challenge is the growing size of the design variable, i.e., the precoding matrix, which leads to growing computational complexity. For problems considering finite constellations, the even more fatal issue is the exponential growth in the cardinality of observed constellations at each user. It can be observed from (5) and (6) that the computational complexity grows exponentially with respect to the number of received signal sources. An estimate of computational complexity of the sum-rate of a MU-MIMO system with different number of users and constellation cardinality is given in Table II in terms of floating point operations (FLOPs). Practically, evaluating scenarios with more than 12 bits transmitted over all the streams becomes difficult.
@@
2 $1.18\times 10^{8}$ $2.77\times 10^{10}$ $7.05\times 10^{12}$
@@
Figure 2: Illustration of user grouping.
To enable precoding designs and evaluation for large-scale systems (e.g., massive MIMO) with practical computational burden and acceptable reduction in performance, we introduce user grouping and null space projection techniques. The idea is to partition the co-scheduled users into groups, enable one common stream for each group, and constrain the precoders such that inter-group interference is eliminated. Figure 2 illustrates a system in which $K$ users (assuming that $K$ is even) are partitioned into $\frac{K}{2}$ groups, and each group contains two users. The elimination of inter-group interference is illustrated since signals carrying the common and private streams intended for a particular user do not propagate to other groups, although intra-group interference, i.e., interference caused by the private stream intended for a co-group user, potentially exists. In contrast to Table II, an estimate of computational complexity with user grouping and interference nulling is given by Table II, where the growth in FLOPs is roughly linear w.r.t. the number of users.
Problem formulation involving user grouping and interference nulling can be written as the following:
$\mathcal{P}_{7}:\;\;\underset{\mathbf{P},\mathbf{c},\mathcal{G}}{\max}\>\>$ $\Omega(\mathbf{P},\mathbf{c},\mathcal{G})$
s.t. $\|\mathbf{P}\|_{\text{F}}^{2}\leq P_{\text{T}}$ (32a)
$|h_{m}^{H}\mathbf{p}_{k}|^{2}=0,$
$\forall m\in\mathcal{K}_{a},k\in\mathcal{K}_{b},\mathcal{K}_{a}\in\mathcal{G},\mathcal{K}_{b}\in\mathcal{G},\mathcal{K}_{a}\neq\mathcal{K}_{b}$ (32b)
$|h_{m}^{H}\mathbf{p}_{\mathcal{K}_{1}}|^{2}=0,\forall m\notin{\mathcal{K}_{a}}\in\mathcal{G}$ (32c)
$\sum_{k\in\mathcal{K}_{a}}C_{k}=R_{\mathcal{K}_{1}},\forall\mathcal{K}_{a}\in\mathcal{G}$ (32d)
$\mathbf{c}\succcurlyeq\mathbf{0}^{1\times K},$ (32e)
where
$\Omega(\mathbf{P},\mathbf{c},\mathcal{G})=\sum_{k\in\mathcal{K}}u_{k}R_{k}^{\text{SIC/SIC-free}}$ (33)
for the WSR problem, and
$\Omega(\mathbf{P},\mathbf{c},\mathcal{G})=\min_{k\in\mathcal{K}}R_{k}^{\text{SIC/SIC-free}}$ (34)
for the MMF problem. $\mathcal{G}=\{\mathcal{K}_{1},\mathcal{K}_{2},...\}$ denotes the grouping, whose elements are sets of user indices allocated to different groups. $\mathbf{P}=[\mathbf{p}_{\mathcal{K}_{1}},\mathbf{p}_{\mathcal{K}_{2}},...,\mathbf{p}_{1},\mathbf{p}_{2},...,\mathbf{p}_{K}]$, where $\mathbf{p}_{\mathcal{K}_{i}}$, $\forall\mathcal{K}_{i}\in\mathcal{G}$, is the common stream precoder intended for users in group $\mathcal{K}_{i}$. $R_{\mathcal{K}_{i}}$, $\forall\mathcal{K}_{i}\in\mathcal{G}$, denotes the achievable rate of the common stream intended for group $\mathcal{K}_{i}$. In addition to the constraints in $\mathcal{P}_{1}$ and $\mathcal{P}_{3}$, (32b) and (32c) are added to $\mathcal{P}_{7}$ to enforce inter-group interference nulling. As a result of (32b) and (32c), computation of the achievable rate of user-$k$, $\forall k\in\mathcal{K}$, can be simplified by only considering signals intended for users in the same group as user-$k$. Specifically, $\forall k\in\mathcal{K}_{i}$, $\forall\mathcal{K}_{i}\in\mathcal{G}$, the computation of (III-B-III-B) and (III-B-III-B) can be reduced by substituting $\mathcal{K}_{i}$ for $\mathcal{K}$ without affecting the results.
$\mathcal{P}_{7}$ involves designing $\mathcal{G}$, i.e., the user grouping, and therefore is a combinatorial optimization problem. The joint optimization of $\mathbf{P}$, $\mathbf{c}$ and $\mathcal{G}$ is difficult. We therefore design $\mathcal{G}$ through a heuristic approach and ensure (32b) and (32c) through multi-stage precoding, followed by optimizing $\mathbf{P}$ and $\mathbf{c}$. To address the growth in the dimension of $\mathbf{P}$ w.r.t. $n_{\text{T}}$, we also utilize a technique to reduce the dimensionality of the design variables to further accelerate optimization algorithms.
@@ V-A User Grouping
It is intuitive that, with an optimal precoding matrix, a user would suffer less interference from streams intended for co-scheduled users which have dissimilar channels. Hence, we heuristically group users with similar channels to limit the potential performance degradation. For the purpose of limiting the computational burden, the number of users within each group should be limited. We consider a special case where each group contains two users. The proposed user grouping algorithm is described in Algorithm 5777We assume $K$ to be even for clarity, while extension to odd numbers is straightforward. For example, one can assume that one group contains only one user, while all other groups contain two users., where $q(\mathbf{h}_{m},\mathbf{h}_{n})$ is a predefined channel similarity metric. We use the cosine similarity of channels for our analysis, i.e.,
@@
although alternative metrics might provide superior results or serve different purposes.
@@
2:  Compute $\mathbf{Q}\in\mathbb{C}^{K\times K}$, where$[\mathbf{Q}]_{m,n}=q(\mathbf{h}_{m},\mathbf{h}_{n})$, $\forall m<n,m\in\mathcal{K},n\in\mathcal{K}$, and$[\mathbf{Q}]_{m,n}=-1$, $\forall m\geq n,m\in\mathcal{K},n\in\mathcal{K}$.
@@
It is evident that Algorithm 5 can be extended to handle cases with more than two users per group.
@@ V-B Null Space Projection and Dimensionality Reduction
$\mathbf{p}_{\mathcal{K}_{i}}=\mathbf{F}_{\mathcal{K}_{i}}\mathbf{G}_{\mathcal{K}_{i}}\mathbf{v}_{\mathcal{K}_{i}},$ (36)
@@
where $\mathbf{F}_{\mathcal{K}_{i}}$ projects the subsequent vector onto the null space of the channels of users outside $\mathcal{K}_{i}$, such that inter-group interference is eliminated. We denote and express the inter-group channels to users in $\mathcal{K}_{i}$ as
@@
where the columns of $\mathbf{H}_{\mathcal{K}_{j}}$ are the channel vectors of users in group $\mathcal{K}_{j}$. With eigenvalue decomposition (EVD), it is readily seen that inter-group interference nulling can be achieved by having
@@
Apart from interference nulling, an additional benefit of $\mathbf{F}_{\mathcal{K}_{i}}$ is that the dimensionality of the subsequent design variable is reduced from $N_{\text{T}}$ to $N_{\text{T}}-K+2$, i.e., the dimensionality of the null space of inter-group channels. We introduce $\mathbf{G}_{\mathcal{K}_{i}}$ to further reduce the dimensionality of the design variable by extending the low-dimensional subspace property (LSP) introduced by [43] to finite constellation scenarios as follows.
@@
With inter-group interference nulling, the LSP implies that optimal $\mathbf{G}_{\mathcal{K}_{i}}\mathbf{v}_{\mathcal{K}_{i}}$ and $\mathbf{G}_{\mathcal{K}_{i}}\mathbf{v}_{k}$ lies in the column space of $\mathbf{F}_{\mathcal{K}_{i}}^{H}\mathbf{H}_{\mathcal{K}_{j}}$. This allows a reduction in the dimensionality in the design variable from $N_{\text{T}}-K+2$ to $2$ while ensuring that the optimal solution can still be achieved. Let
$\mathbf{F}_{\mathcal{K}_{i}}^{H}\mathbf{H}_{\mathcal{K}_{i}}=\mathbf{U}_{\mathcal{K}_{i}}\mathbf{\Sigma}\mathbf{V}_{\mathcal{K}_{i}}^{H}$ (40)
be compact singular value decomposition (SVD) of $\mathbf{F}_{\mathcal{K}_{i}}^{H}\mathbf{H}_{\mathcal{K}_{j}}$. We restrict $\mathbf{G}_{\mathcal{K}_{i}}\mathbf{v}_{k}$ to be in the desired subspace by choosing [^8] To achieve the same purpose, one can also have $\mathbf{G}_{\mathcal{K}_{i}}=\mathbf{F}_{\mathcal{K}_{i}}^{H}\mathbf{H}_{\mathcal{K}_{j}}$. However, it was observed numerically that the proposed $\mathbf{G}_{\mathcal{K}_{i}}$ leads to better convergence when optimizing $\mathbf{v}_{k}$. One potential reason is that the proposed $\mathbf{G}_{\mathcal{K}_{i}}$ guarantees that $\|\mathbf{p}_{k}\|^{2}=\|\mathbf{v}_{k}\|^{2}$ and $\|\mathbf{p}_{\mathcal{K}_{i},\text{c}}\|^{2}=\|\mathbf{v}_{\mathcal{K}_{i}}\|^{2}$.
@@
The procedures of interference nulling and dimensionality reduction are summarized in Algorithm 6. Given solutions for the first and second stages of precoding, the third stage, i.e., $\mathbf{v}_{k}$ and $\mathbf{v}_{\mathcal{K}_{i}}$, is designed in a manner similar to that in Section IV. Note that, as a result of inter-group interference nulling, inter-group interference does not need to be considered when computing the achievable rates and their gradients for $\mathcal{P}_{1}$ and $\mathcal{P}_{3}$. For the MMF problem, as each group posses a common stream, the design of common rate allocation, i.e., Algorithm 3, is carried out for each group separately.
Algorithm 6 Interference Nulling and Dimensionality Reduction
0:  $\mathcal{G}$, $\mathbf{h}_{k}$, $k\in\mathcal{K}$.
0:  $\mathbf{F}_{\mathcal{K}_{i}}$ and $\mathbf{G}_{\mathcal{K}_{i}}$, $\forall\mathcal{K}_{i}\in\mathcal{G}$.
1:  Construct $\mathbf{H}_{\bar{\mathcal{K}}_{i}}$, $\forall\mathcal{K}_{i}\in\mathcal{G}$.
2:  Perform EVD to $\mathbf{H}_{\bar{\mathcal{K}}_{i}}\mathbf{H}_{\bar{\mathcal{K}}_{i}}^{H}$ to obtain $\tilde{\mathbf{V}}_{\bar{\mathcal{K}}_{i}}$, $\forall\mathcal{K}_{i}\in\mathcal{G}$.
3:  $\mathbf{F}_{\mathcal{K}_{i}}:=\tilde{\mathbf{V}}_{\bar{\mathcal{K}}_{i}}$, $\forall\mathcal{K}_{i}\in\mathcal{G}$.
4:  Perform SVD to $\mathbf{F}_{\mathcal{K}_{i}}^{H}\mathbf{H}_{\mathcal{K}_{i}}$ to obtain $\mathbf{U}_{\mathcal{K}_{i}}$ and $\mathbf{V}_{\mathcal{K}_{i}}$, $\forall\mathcal{K}_{i}\in\mathcal{G}$.
5:  $\mathbf{G}_{\mathcal{K}_{i}}:=\mathbf{U}_{\mathcal{K}_{i}}\mathbf{I}\mathbf{V}_{\mathcal{K}_{i}}^{H}$.
6:  return  $\mathbf{F}_{\mathcal{K}_{i}}$ and $\mathbf{G}_{\mathcal{K}_{i}}$, $\forall\mathcal{K}_{i}\in\mathcal{G}$.
V-C Complexity Analysis
For clarity, we assume that all private streams apply the same alphabet, denoted by $\mathcal{X}$. For the general approaches proposed in Section IV, computing (III-B), (III-B) and their gradients w.r.t. $\mathbf{P}$ requires $\mathcal{O}(|\mathcal{X}_{\text{c}}|^{2}|\mathcal{X}|^{2K}Kn_{\text{T}})$ operations. Computing (III-B) and its gradient requires $\mathcal{O}(|\mathcal{X}|^{2K}Kn_{\text{T}})$. Therefore, the complexity of Algorithms 1 and 4 is $\mathcal{O}(I_{\text{it}}|\mathcal{X}_{\text{c}}|^{2}|\mathcal{X}|^{2K}K^{2}n_{\text{T}})$, where $I_{\text{it}}$ is the number of iterations.
With user grouping, interference nulling and dimensionality reduction, $\mathcal{O}(|\mathcal{X}_{\text{c}}|^{2}|\mathcal{X}|^{2})$ is required for (III-B), (III-B) and their gradients. $\mathcal{O}(|\mathcal{X}|^{2})$ is required for (III-B) and its gradient. For user grouping, Algorithm 5 with the proposed similarity metric requires $\mathcal{O}(K^{2}N_{\text{T}})$. The complexity of EVD for an $n\times n$ matrix is $\mathcal{O}(n^{3})$, and that of SVD for an $m\times n$ matrix is $\mathcal{O}(\min(mn^{2},m^{2}n))$. Hence, Algorithm 6 requires $\mathcal{O}(n_{\text{T}}^{3}+n_{\text{T}}^{2}K+\min(n_{\text{T}}K^{2},n_{\text{T}}^{2}K))$. The overall complexity of user grouping, interference nulling, dimensionality reduction, and WSR/MMF optimization proposed in this section requires $\mathcal{O}(I_{\text{it}}|\mathcal{X}_{\text{c}}|^{2}|\mathcal{X}|^{2}K+n_{\text{T}}^{3}+n_{\text{T}}^{2}K+n_{\text{T}}K^{2})$. Compared to $\mathcal{O}(I_{\text{it}}|\mathcal{X}_{\text{c}}|^{2}|\mathcal{X}|^{2K}Kn_{\text{T}})$ for the general algorithm, the prohibitive exponent $K$ disappears, although there is an increase in the power of $n_{\text{T}}$ and $K$ due to EVD and SVD.
@@ VI-A Channel Model
where $\mathbf{h}_{k}^{\text{LoS}}(\theta_{k}^{\text{az}},\theta_{k}^{\text{el}})$, $\mathbf{h}_{k}^{\text{NLoS}}$ and $\kappa_{k}$ denote the Line-of-Sight (LoS) channel component, non-LoS channel component and and the Rician factor. $\mathbf{h}_{k}^{\text{LoS}}$ is characterized by the array geometry and the direction of departure in azimuth and elevation, denoted by $\theta_{k}^{\text{az}}$ and $\theta_{k}^{\text{el}}$ respectively. $\mathbf{h}_{k}^{\text{NLoS}}$ is modeled by Rayleigh fading with all the entries follow $\mathcal{CN}(0,1)$.
In this work, we aim to model channels in practical scenarios where users are in high-density demand areas, therefore simulating users in similar directions. For small-scale systems, a uniform linear array is assumed along the $y$ axis with $\theta_{k}^{\text{el}}=0,\forall k\in\mathcal{K}$ since elevation does not affect the channels. For large-scale systems, a uniform rectangular array along the $y$-$z$ plane is assumed. It is assumed that $\kappa_{1}=...=\kappa_{K}=\kappa$.
@@ VI-B Adaptive Modulation
The advantage of RSMA lies in its flexibility in adjusting the power and rates between common and private streams. To fully exploit this advantage under finite constellations, we allow RSMA to dynamically adjust the constellations. In particular, we enable RSMA to maximize the objective function by exhaustively searching for constellations from a pre-defined dictionary of transmission modes. To reduce the searching complexity, private streams are limited to using the same constellation in each transmission mode. To ensure fairness, different modes within the same dictionary are designed to produce the same maximum achievable rate, denoted by $R_{\text{max}}$. According to the above design principles, the four transmission mode dictionaries used in the simulations are documented in Table VI - VI in Appendix C. Note that with Mode 1 in all the examples, RSMA is equivalent to SDMA, since the precoding optimization will allocate zero power to the common stream when the common stream constellation is empty. For large-scale systems, we assume that all the groups use the same transmission mode, which is searched from one of the pre-defined dictionaries. In the numerical results, where necessary, $K$ and $R_{\text{max}}$ are provided in the legends or captions to identify the transmission mode dictionary in use.